OpenStudy (anonymous):

y= -2+2i and z=1+i(squareroot)3 find yz in polar form

OpenStudy (anonymous):

you can put them in polar form first, then multiply

OpenStudy (anonymous):

do you know how to put them in polar form?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

you need two numbers, \(r=\sqrt{a^2+b^2}\) and \(\theta\) where \(\tan(\theta)=\frac{b}{a}\)

OpenStudy (anonymous):

for \(-2+2i\) you have \(a=-2,b=2\) and so \(r=\sqrt{2^2+2^2}=\sqrt{8}=2\sqrt{2}\)

OpenStudy (anonymous):

then \(\tan(\theta)=\frac{2}{-2}=-1\) and so because you are in quadrant 2, \(\theta =\frac{3\pi}{4}\)

OpenStudy (anonymous):

oops typo there \[-2+2i=2\sqrt{2}\left(\cos(\frac{3\pi}{3})+i\sin(\frac{3\pi}{4})\right)\]

OpenStudy (anonymous):

i was wondering where that came from

OpenStudy (anonymous):

yeah it was a mistake

OpenStudy (anonymous):

now repeat the process for \(1+\sqrt{3}i\)

OpenStudy (anonymous):

in this case \(a=1,b=\sqrt3\) and so \[r=\sqrt{1^2+\sqrt{3}^2}=\sqrt{4}=2\]

OpenStudy (anonymous):

also \(\tan(\theta)=\sqrt{3}\) and since you are in quadrant 1 you have \(\theta=\frac{\pi}{3}\)

OpenStudy (anonymous):

gives \[1+\sqrt{3}i=2\left(\cos(\frac{\pi}{3})+i\sin(\frac{\pi}{3})\right)\]

OpenStudy (anonymous):

the last part is easiest to multiply the two numbers, add the angles and multiply the modulus (absolute value, \(r\))

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

yw